# Leveraging Generative AI for Smart City Digital Twins: A Survey on the Autonomous Generation of Data, Scenarios, 3D City Models, and Urban Designs

### v1

@@ 3.2.4 Urban Disaster Management
As detailed examples, Hofmann and Schüttrumpf (2021) presents a novel deep learning approach for real-time pluvial flood prediction. The study introduces floodGAN, a deep convolutional generative adversarial network, which significantly accelerates flood prediction while maintaining high accuracy. This approach effectively utilizes the capabilities of deep learning, particularly GANs, to understand and predict complex flood dynamics based on spatially distributed rainfall events. The innovation lies in its ability to provide fast, accurate flood predictions essential for real-time applications such as early warning systems, marking a substantial advancement in flood modeling and risk management. Salih et al. (2022) explores the use of GANs and Mask Region-Based Convolutional Neural Networks (Mask-R-CNN) for floodwater detection and segmentation in images. The study highlights the capability of GANs to generate segmented images of flooded areas, emphasizing their potential in applications where quick and accurate floodwater identification is crucial. The research demonstrates the effectiveness of GANs in image-based flood detection, offering a novel approach to enhancing flood response and management strategies. Yadav et al. (2024) introduces an innovative unsupervised approach for flood mapping using Synthetic Aperture Radar (SAR) time series data. It leverages a Variational Autoencoder (VAE) trained with reconstruction and contrastive learning techniques. This model is adept at detecting changes by examining the latent feature distributions between pre-flood and post-flood data. Its performance, evaluated against nine different flood events, demonstrates superior accuracy over existing unsupervised models. This approach marks a significant advancement in the application of deep learning to environmental monitoring, specifically in enhancing flood detection capabilities using satellite data. Lago et al. (2023) introduces cGAN-Flood, a conditional generative adversarial network developed to enhance flood prediction in urban catchments. Unlike traditional ANN models, cGAN-Flood can generalize flood predictions to areas outside its training dataset, addressing varying boundary conditions. It achieves this by using two cGAN models for identifying wet cells and estimating water depths, trained with HEC-RAS outputs. The method was tested on different urban catchments and compared with HEC-RAS and the rapid flood model WCA2D. It successfully predicted water depths, despite some underestimation in channels during low-intensity events. Significantly faster than existing models, cGAN-Flood offers a promising, efficient alternative for large-scale flood forecasting. Delacruz (2020) presents an innovative approach using GANs for rapid classification of earthquake-induced damage to railways and roads. It highlights the use of GANs to generate synthetic images of damage, addressing the challenge of limited real-world data for training deep learning models. This method demonstrates the potential of GANs in enhancing the efficiency and accuracy of structural damage assessment, which is crucial for emergency response and infrastructure recovery efforts. Zhou et al. (2023a) utilizes GANs for urban morphology prediction related to local climate zones (LCZs). It employs GANs to integrate LCZ and urban morphology data from six cities, demonstrating superior accuracy and convergence capabilities, particularly with the Pix2pix model. This approach allows for rapid and responsive 3D urban morphology prediction, offering significant insights for enhancing local climates in urban areas through 3D generative design techniques.
3.2.5 Urban Water Management
Currently, the applications of generative AI in urban water management is relatively limited, especially when compared to its broader use in other urban sectors. The primary focus of existing applications lies in the prediction and structural management of sewer water systems. This section will selectively explore and discuss a handful of pertinent examples in this domain. Koochali et al. explores the innovative usages of GANs in urban water management, specifically in improving combined sewer overflow predictions, such as depth and flow rate. GANs are used for data augmentation, generating synthetic time series to balance datasets, thus enhancing the accuracy of prediction models. The study demonstrates that models trained with synthetic data generated by GANs can perform comparably to those trained with real data. The study proposed novel deep-learning approaches that can address challenges like data scarcity in urban water management. Fu et al. (2022) focuses on the application of GANs for tasks like predicting water system behaviors, improving operational efficiency, and enhancing decision-making processes. The paper underlines the effectiveness of GANs in handling complex, data-rich scenarios typical in urban water systems, demonstrating how these advanced AI models can contribute to more sustainable and effective water management strategies. Situ et al. (2021) presents a novel approach for sewer defects detection. It utilizes state-of-the-art Style-based Generative Adversarial Networks (StyleGANs) to generate high-quality synthetic images of sewer defects. These images are then used to train and fine-tune seven well-known CNN models for automated sewer defects detection. The study demonstrates that both variants of StyleGANs are efficient in producing diverse and detailed images, significantly enhancing the performance of deep learning models in sewer defects detection.
3.3 3D City Modeling
The rapid advancement of City Digital Twins is propelled by digital infrastructure and ICT technologies linked to the Internet of Things (IoT), leading to the transformation of traditional cities into smart cities with enhanced infrastructure maintenance, socio-economic growth, and urban governance (Wang et al., 2021; Caprari et al., 2022). Among the majority of the existing digital urban twins, 3D city models has played a pivotal role to support the 3D modelling, monitoring and visualization of the smart city (Masoumi et al., 2023). 3D city models are also developed to enable the interactive and immersive visualization of complex urban environments to facilitate public engagement and collaborative urban planning (Dembski et al., 2019; Xu et al., 2023a), simulation and evaulation of smart technologies (Xu et al., 2023c, d), and 3D cadastral and urban design (Shojaei et al., 2013; Polys et al., 2018).
Despite that the traditional geomatics and procedure-based methods (Lerma et al., 2010; Bejleri et al., 2021) are capable of creating the Digital Surface Models (DSMs) of many large cities (Singh et al., 2013), these DSMs have limited resolution (e.g., texture and geometry) for representing building and infrastructure at the acceptable Level of Details (LoD). Additionally, many novel uses of 3D digital models (e.g., Hardware-In-the-Loop (HIL) simulations and scientific gamification) would require a gamified environment in which urban entities, such as buildings, roads, and infrastructure, are fully separated and have individual physical and gaming properties (Xu et al., 2023c, a). In this setting, creating a high-fidelity 3D representation of a city is a challenging effort that often requires massive, repetitive, time-consuming and manual efforts for surveying and 3D modeling of urban entities (Xu et al., 2023a), along with advanced computing resources for processing and integrating large-scale GIS, remote-sensing (e.g., photogrammetry and LIDAR scan), and crowd-sourcing data in a virtual environment (Xu et al., 2023c). It has been an emerging trend to leverage the 3D content-generation capability of the recent generative AI techniques to automate the tremendous manual efforts for creating 3D assets of urban entities, such as terrain, buildings, infrastructures, and vegetation (Xu et al., 2023c).
3.4 3D Building Generation
The autonomous capability of recent generative AI models has allowed them to create photo-realistic 3D representations of buildings and street environments(Pang and Biljecki, 2022; Wu et al., 2022). These applications have advanced the technical capability of many urban science fields, including architecture design, building construction and operations, and urban planning. Du et al. (2020b) introduces a novel method for 3D building fabrication using a Hybrid GAN. This method incorporates a unique loss function, including cycle consistency and perceptual loss, and utilizes a multi-properties GAN chain to create complex architectural designs. It also features a mixed GAN network for generating geometry and texture coordination. This approach significantly enhances the realism of 3D building models, achieving a 20% performance improvement over traditional methods, demonstrating its potential in applications like digital city construction and smart urban planning. (Kelly et al., 2018) discusses an innovative framework called FrankenGAN. This system is designed to enhance coarse building mass models by adding detailed geometric and texture details, creating realistic and visually appealing urban landscapes. FrankenGAN employs a cascade of Generative Adversarial Networks (GANs), synchronized to maintain consistent styles across multiple scales in neighborhoods. It offers user control over the variability of output styles through exemplar images, allowing for a blend of automated and user-guided design. This technology is significant for its potential to revolutionize urban modeling and architectural design, providing a powerful tool for researchers and professionals in these fields to generate detailed, large-scale 3D city models efficiently and with high fidelity. Bacharidis et al. (2020) explores the use of GANs for reconstructing 3D building facades from images. It employs deep learning techniques for facade structural element detection, depth point-cloud generation, and protrusion estimation. The study demonstrates GANs’ capability in segmenting building components from images and generating detailed 3D facades. These capabilities show GANs’ potential in evaluating urban scenes and extracting urban design parameters, such as window ratio, skyline, and other applications, including cultural heritage preservation. The approach aims to address the challenges of architectural diversity and data scarcity in 3D modeling.
3.5 3D Street Environment Construction
A few other studies have developed generative AI strategy to digitize the entire street environment (with buildings and infrastructure) as 3D models, laying an foundation for the next-generation extended reality applications for supporting participatory and collaborator urban design and planning (Xu et al., 2023a). Tancik et al. (2022) presents a significant advancement in neural rendering for large-scale urban environments. It introduces Block-NeRF, a variant of Neural Radiance Fields (NeRF), specifically designed to handle city-scale scenes. This method effectively decomposes a large scene into individually trained NeRFs, allowing for scalable rendering and enabling updates to specific parts of the environment without retraining the entire network. This technique represents a major step forward in the field of photorealistic 3D reconstruction and neural rendering. (Kim et al., 2020b) introduces a novel method for generating three-dimensional virtual models of cities from a single street-view image. This approach differs from traditional reconstruction methods by utilizing machine learning, specifically generative adversarial networks (GANs) and convolutional neural networks (CNNs), to create terrain maps and identify city components and styles. The generated 3D virtual cities are visually plausible and resemble actual cities. This method represents a significant advancement in the field of urban modeling, offering a unique and efficient way to create detailed, large-scale city models, which is particularly useful for applications in gaming, film, and urban planning. The use of GANs and CNNs in this context is notable for its ability to support research and development in these areas by providing a more streamlined and user-friendly approach to city modeling.
3.6 Generative Urban Design and Optimization
Generative design is defined as using computer-aided generative methods to produce design options, practices, and solutions, improving certain stages of the typical design process (Stojanovski et al., 2020; BuHamdan et al., 2021). In recent years, generative AI models play a crucial role as the intelligent agent in computer-aided urban design. They are employed in many studies of urban design alternatives and strategies that are driven by urban big data and previous solutions (Al Qeisi and Al-Alwan, 2021). The trend of publications that apply generative AI in urban design and optimization has been summarized in Figure 3. Recent progress in generative urban design is evident in the urban science and smart city sector, where these models are increasingly used for tasks such as urban layout and land use optimization, environmental impact assessments, and urban resource management. At a high level, a study by Quan et al. (2019) highlighted the effectiveness of AI in managing urban complexity and sustainability, as well as supporting generative and participatory urban design. Furthermore, the integration of AI in smart cities is demonstrated by the work of Batty (2018), who emphasized AI’s role in analyzing large-scale urban data to inform better city planning and management. The trend towards AI-aided urban planning signifies a paradigm shift, promising smarter, more efficient, and sustainable urban environments.
Figure 3: Summary of generative urban design and optimization publications using Scopus database (accessed on Jan 10th 2024)
As examples of detailed generative urban design applications, Huang et al. (2022) explores the use of GANs, in urban design for environmental performance optimization. It demonstrates the use of GANs in rapidly generating urban design alternatives that meet specific environmental criteria, highlighting deep learning’s ability to process complex spatial data and simulate various urban development scenarios. This approach represents a significant innovation in integrating environmental considerations into the urban design process using advanced AI techniques. Fedorova (2021) examines the use of deep learning, particularly GANs, in creating urban layout designs. It focuses on GANs’ ability to generate diverse and realistic urban layouts, which can significantly aid in urban planning and development. This research underscores the potential of deep learning in addressing complex spatial design challenges, providing innovative tools for urban designers and planners. Noyman and Larson (2020) explores the application of generative AI models, such as GANs and VAEs, in urban design, focusing on generating urban block layouts. This innovative approach does not require predefined parameters for design solutions, as the neural network learns the visual features of existing city structures. The research demonstrates GANs’ ability to adapt designs to various urban contexts and morphologies. Significantly, it highlights the potential of generative AI in urban planning, offering a novel method for designing urban blocks that structurally align with the existing urban fabric without the need for explicit parameter definitions. It should be mentioned that social, economic, and environmental considerations can be used as constraints to improve the urban blocks from the livability and sustainability perspectives. Quan (2022) develops Urban-GAN, an AI-aided generative urban design computation system, which empowers public participation in urban design through Generative Adversarial Networks (GANs) and Deep Convolutional Neural Networks (DCNNs). This innovative system enables individuals with little design expertise to generate professional-level urban designs. It employs sub-symbolic representation and case-based reasoning, facilitating users to select urban form cases and generate similar design schemes. The system’s application in hypothetical design experiments across cities like Manhattan, Portland, and Shanghai illustrates its effectiveness in generating diverse urban forms, highlighting its potential in democratizing urban design and planning. KIM et al. (2020) focuses on developing an AI tool called PlacemakingAI, which uses Generative Adversarial Networks (GANs) for urban design. The tool facilitates public participation in urban planning, allowing users to visualize sustainable urban spaces in real-time. It leverages both supervised and unsupervised GANs to generate synthetic images from datasets of walkable streets. This research is significant for its innovative use of GANs in urban design, enabling a participatory approach and fostering engagement from citizens, designers, and stakeholders in the urban design process.
4 Future of Generative AI in Urban Digital Twins
@@
This section explores the integration of generative AI models with emerging urban digital twin applications (e.g., web-based platforms, cyber-physical systems, and cyberinfrastructure), aiming to augment their current capabilities for more efficient smart city management. Additionally, we delve into the potential technical and methodological challenges associated with the development of generative urban digital twins, underscoring the complexities and innovations required for this visionary approach.
4.1 Cognitive Digital Twins: Opportunities
As an integration between AI and digital twins, cognitive digital twins represent a significant advancement in the digital twin concept, integrating enhanced capabilities beyond their traditional counterparts. These advanced models assimilate cognitive functions, enabling them to dynamically process and learn from data. They harness deep learning to analyze complex datasets, predict future scenarios, and make decisions autonomously, leading to improved efficiency and adaptability. In urban science and smart city applications, cognitive digital twins are proving invaluable for enabling the real-time situational awareness and predictive analytics of complex urban systems using a variety of neuron networks for classification, faulty detection, object detection, and optimization purposes. They also facilitate human-AI collaboration by involving users (e.g., experts, policymakers, and city residents) in the urban management processes through visual dashboards and visualizations. Nevertheless, cognitive digital twin applications that leverage the capability of generative AI to solve complex urban problems are currently limited in the smart city sector.
@@ 4.1 Cognitive Digital Twins: Opportunities
Figure 4: Potential opportunities of generative AI techniques for enhancing cognitive urban digital twins at the conceptual level.
The vision of incorporating generative AI techniques in cognitive urban digital twins is to achieve an autonomous and generative system that will be cognitive, context-aware, fully automated, and cost-effective. This paper believes that the capability of previous generative AI applications reviewed in the previous section can be seamlessly integrated into cognitive urban digital twins. Future directions or opportunities for the generative urban digital twins are discussed as follows (as depicted in Figure 4).
Enhancing Situational Awareness:
Through data augmentation, generative AI techniques can generate realistic urban data to refine the spatiotemporal resolutions and facilitate anomaly detection in the raw data, such as point clouds and satellite images (Wu et al., 2022). Therefore, these techniques enhance the urban digital twins’ capability to replicate the dynamics and reflect the state of the real-world urban systems in the digital environment.
Improving Prediction Accuracy:
@@
With improved data quantity, quality, and realistic synthetic data that are generated based on real-world data, the urban digital twins’ predictive analytics capability can be improved to predict or forecast the dynamics and response of different urban systems to either futuristic or hypothetical scenarios through data and physics-driven models. For instance, the design and development of critical infrastructure such as airports and their complex interactions with urban transportation systems can be modeled and simulated through different scenarios by integrating proposed future developments (e.g., buildings and transport networks) with current and real-time data (e.g.traffic, current buildings, and land uses). These ”what-if” scenarios also provide alternatives to digital twin users for more adaptive decision support.
Facilitating Human-AI Collaboration:
By generating visualizations, text, and data-driven insights that are understandable to humans, generative AI models can facilitate better collaboration between humans and AI systems. These models can be designed to learn more interpretable features, which are more aligned with human-understandable concepts derived from the domain knowledge. The explanations derived from these models can be made more accessible and sensible to digital twin users.
Enabling Autonomous System Development:
The ability of generative AI models to generate realistic data and codes in different programming languages can be applied to automate the software development of urban digital twins’ components, such as 3D city models and scenario-based simulations (e.g., traffic, mobility, and building energy). This generative ability can significantly increase the adaptability and scalability of digital twin platforms by extending their coverage to new geographic areas and urban domains. For instance, the process of building development approvals, which involves structural, energy, and site examinations, can be automated.
Supporting Decision Optimization:
Generative AI models can generate counterfactuals based on multiple ”what-if” scenarios. These generated urban design or planning strategies can be used as alternatives to help urban planners evaluate the feasibility and efficiency of numerous planning or design alternatives and select the optimized solutions. This addresses the current gap in comprehensive impact analysis of urban design scenarios. This is due to the limitation of computational models and algorithms. For instance, in a Transit Oriented Development (TOD) scenario, different urban design analytics, including building height, sky view factor, walkability, and accessibility to different modes of transport, need to be evaluated.
Promoting AI-aided Design:
Generative models can rapidly generate a wide range of urban design concepts based on specific criteria or policy (e.g., walkability, accessibility, safety, and security, climate resilience), helping planners explore creative ideas and inspirations that might not be immediately visible to humans. Additionally, these models can tailor and modify existing human designs to meet a community’s or environment’s specific needs, and suggest design elements that are more likely to be successful in a particular area.
4.2 Generative AI-Digital Twin Integration

### v2

@@ 3.5.5 Disaster Management
As detailed examples, Hofmann and Schüttrumpf (2021) presents a novel deep learning approach for real-time pluvial flood prediction. The study introduces floodGAN, a deep convolutional generative adversarial network, which significantly accelerates flood prediction while maintaining high accuracy. This approach effectively utilizes the capabilities of deep learning, particularly GANs, to understand and predict complex flood dynamics based on spatially distributed rainfall events. The innovation lies in its ability to provide fast, accurate flood predictions essential for real-time applications such as early warning systems, marking a substantial advancement in flood modeling and risk management. Salih et al. (2022) explores the use of GANs and Mask Region-Based Convolutional Neural Networks (Mask-R-CNN) for floodwater detection and segmentation in images. The study highlights the capability of GANs to generate segmented images of flooded areas, emphasizing their potential in applications where quick and accurate floodwater identification is crucial. The research demonstrates the effectiveness of GANs in image-based flood detection, offering a novel approach to enhancing flood response and management strategies. Yadav et al. (2024) introduces an innovative unsupervised approach for flood mapping using Synthetic Aperture Radar (SAR) time series data. It leverages a VAE trained with reconstruction and contrastive learning techniques. This model is adept at detecting changes by examining the latent feature distributions between pre-flood and post-flood data. Its performance, evaluated against nine different flood events, demonstrates superior accuracy over existing unsupervised models. This approach marks a significant advancement in the application of deep learning to environmental monitoring, specifically in enhancing flood detection capabilities using satellite data. Lago et al. (2023) introduces cGAN-Flood, a conditional generative adversarial network developed to enhance flood prediction in urban catchments. Unlike traditional ANN models, cGAN-Flood can generalize flood predictions to areas outside its training dataset, addressing varying boundary conditions. It achieves this by using two cGAN models for identifying wet cells and estimating water depths, trained with HEC-RAS outputs. The method was tested on different urban catchments and compared with HEC-RAS and the rapid flood model WCA2D. It successfully predicted water depths, despite some underestimation in channels during low-intensity events. Significantly faster than existing models, cGAN-Flood offers a promising, efficient alternative for large-scale flood forecasting. Delacruz (2020) presents an innovative approach using GANs for rapid classification of earthquake-induced damage to railways and roads. It highlights the use of GANs to generate synthetic images of damage, addressing the challenge of limited real-world data for training deep learning models. This method demonstrates the potential of GANs in enhancing the efficiency and accuracy of structural damage assessment, which is crucial for emergency response and infrastructure recovery efforts. Zhou et al. (2023a) utilizes GANs for urban morphology prediction related to local climate zones (LCZs). It employs GANs to integrate LCZ and urban morphology data from six cities, demonstrating superior accuracy and convergence capabilities, particularly with the Pix2pix model. This approach allows for rapid and responsive 3D urban morphology prediction, offering significant insights for enhancing local climates in urban areas through 3D generative design techniques.
4 Future of GenAI in Urban Digital Twins
@@
To address research question Q4, proposed in Section 1.4.2, regarding how GenAI technologies can benefit urban digital twins, this section explores the integration of GenAI models with existing urban digital twin applications (e.g., web-based platforms, cyber-physical systems, cyberinfrastructure, and visual dashboards). The aim is to augment their current capabilities for more efficient participatory smart city management through human-AI collaboration. Additionally, we delve into the potential technical and methodological challenges associated with the development of generative urban digital twins, underscoring the complexities and innovations required for this visionary approach.
@@
Figure 4: Visions and opportunities of GenAI models for enhancing cognitive urban digital twins at the conceptual level.
4.1 Cognitive Digital Twins: Visions and Opportunities
As an integration between AI and digital twins, the emerging Cognitive Digital Twins (CDTs) represent a significant advancement in the digital twin concept, enabling self-learning, memory, and reasoning capabilities beyond their traditional counterparts (Niloofar et al., 2023). These AI models assimilate cognitive functions, enabling CDTs to dynamically process and learn from data. They harness deep learning to analyze complex datasets, predict future scenarios, and make decisions autonomously, leading to improved efficiency and adaptability (Ali et al., 2021). In urban science and smart city applications, the CDT concept illuminates more intelligent, efficient, and automated solutions to complex urban management problems by leveraging the advanced capabilities of various neural networks for classification, fault detection, object detection, and optimization.
Recent revolutionary advancements in GenAI can be integrated into CDT to further enhance its capability for more efficient and participatory urban management in the aspects of (a) facilitating data enhancement and enrichment, (b) building autonomous systems, and (c) promoting human-AI partnership. We present our vision of how GenAI models could be integrated into CDT for smart cities through Figure 4, in which we adopted the core layers of a CDT defined by previous studies (Shi et al., 2022b; Abburu et al., 2020) and elaborated on who individual layers could be benefited from GenAI through the following subsections. The essential layers of a CDT are represented by dark blue boxes, while our proposed visions and opportunities are depicted in green boxes in the figure.
4.1.1 Data Enhancement and Enrichment
Data quality and utility is critical to data-driven applications, especially urban digital twins that leverage urban data to enable the situational awareness, predictive analytics, and informed decision supports of various urban subsystems.
Data Enhancement
refers to the augmentation of existing data sets by adding more detailed information, often from external sources, to make the data more comprehensive. In this context, GenAI models, such as GANs and VAEs, have been widely utilized in contemporary studies to generate realistic data (Wu et al., 2022a), thereby enhancing the reliability and accuracy of a CDT’s situational awareness capabilities. Additionally, VAEs are particularly effective in facilitating fault and anomaly detection, which further supports data enhancement efforts (Guo et al., 2020). VAEs operate by encoding input data into a latent space and then decoding it back to the original space, a process that enables them to capture complex data distributions and identify deviations from normal patterns. This ability to detect anomalies and generate synthetic data contributes to creating more robust and comprehensive datasets, ultimately improving the quality and utility of the data.
Data Enrichment
involves the integration of additional data to provide more context, making the data more meaningful and useful for analysis. In this regard, GenAI models, such as GANs and VAEs, have been frequently utilized in urban studies to generate synthetic data and simulate scenarios (Figueira and Vaz, 2022).
@@ 4.1.1 Data Enhancement and Enrichment
The data synthesis capabilities of these models enhance an urban digital twin’s predictive analytics, allowing for accurate projections and forecasts of urban dynamics in futuristic or hypothetical management scenarios. For instance, the design and development of critical infrastructure, such as airports, and their complex interactions with urban transportation systems can be modeled and simulated through various scenarios. By integrating proposed future developments (e.g., new buildings and transport networks) with current and real-time data (e.g., traffic patterns, existing buildings, and land uses), these models enable the creation of ”what-if” scenarios. These scenarios provide digital twin users with alternatives for more adaptive and informed decision-making, ultimately supporting more resilient and efficient urban planning and management. By leveraging GenAI, urban digital twins can achieve more comprehensive, context-rich data, ultimately leading to more reliable data-driven insights and inferences, as well as more accurate predictive analytics.
This enhancement creates opportunities for making better-informed urban decisions and strategies, providing effective solutions to address digital twin challenges C1 and C2 discussed in Section 1.2.
4.1.2 Privacy-Preserving Synthetic Data Generation
Privacy-Preserving Synthetic Data Generation refers to the use of advanced generative AI techniques, such as GANs and VAEs, to create artificial data that closely mimics real-world data while safeguarding the privacy of the individuals whose information is being represented (Liu et al., 2022; Ghatak and Sakurai, 2022). This approach is particularly crucial in urban research, where the collection and utilization of extensive datasets, such as mobility trajectories and building occupancy data, pose significant privacy and sovereignty concerns. In urban research, mobility trajectory data, which tracks individuals’ movements across a city, can reveal sensitive information about their daily routines, frequently visited locations, and social interactions. Similarly, building occupancy data can disclose patterns related to the presence and activities of individuals within specific buildings, raising potential privacy risks. By generating synthetic data that replicates the statistical properties and patterns of these datasets, researchers can conduct analyses and derive insights without exposing personal information (Abay et al., 2019).
Generative AI models are instrumental in this process due to the advantages provided by their architectures. GANs consist of two neural networks—a generator and a discriminator—that work in tandem to produce data indistinguishable from real data. The generator creates synthetic data, while the discriminator evaluates its authenticity. Through iterative training, GANs improve the quality of the synthetic data, ensuring it accurately reflects real-world patterns. VAEs, on the other hand, encode real data into a latent space and then decode it back into synthetic data, preserving the essential characteristics while obscuring individual-specific details. This approach ensures that the generated data maintains the integrity and variability of the original data while enhancing privacy and reducing the risk of re-identification. By employing these generative AI techniques, urban researchers can create privacy-preserving synthetic datasets that enable robust analysis and innovation. This method not only addresses ethical and legal concerns related to data privacy but also enhances the availability of high-quality data for urban planning, policy-making, and smart city development.
In regards, privacy-preserving synthetic data generation can effectively address the challenges posed by data ownership, data residency, data colonialism by creating artificial datasets that mimic real data without containing any actual sensitive and protected information (Boedihardjo et al., 2022). This approach allows organizations to comply with local data residency laws and facilitates ethical cross-border data sharing, mitigating data colonialism by enabling ethical data sharing (Majeed, 2023; Vie et al., 2022). Additionally, it empowers local stakeholders by enabling them to generate and control representative datasets, thus reducing dependency on external entities and promoting fair data practices ,and providing effective solutions to address digital twin challenge C5 discussed in Section 1.2.
4.1.3 Autonomous Systems
Recent developments in GPT and RAG have highlighted the potential for creating autonomous systems or agents to automate both data analytics and software development (Suri et al., 2023). Building an urban digital twin requires significant interdisciplinary efforts in data management, analysis, and software engineering. In this regard, integrating GenAI models, particularly GPTs, can help automate labor-intensive, time-consuming, and repetitive tasks, thereby reducing the cost of developing and maintaining an urban digital twin. GenAI-powered autonomous systems can be deployed in the virtual and service layers of a CDT to generate computer code in various programming languages, facilitating the creation of core data analytics and software components.
Autonomous Data Analytics
leverage GPTs’ reasoning and generative capabilities to develop automated data workflows and pipelines that can perform advanced data analysis (e.g., statistical, mathematical, and spatial analysis) by calling appropriate libraries based on user-defined prompts (Vemprala et al., 2023). These autonomous systems often employ prompt engineering, predefined prompt bases and templates, and fine-tuning techniques to instruct large language models on specialized data analysis tasks. An example of an autonomous spatial analysis tool powered by the ChatGPT API is presented by Li and Ning (2023). With the recent advancements in RAG, GPTs are now empowered with the capability to query external information sources through vector databases, broadening the application and utility of these generative models.
Autonomous Software Development
focuses on utilizing GPTs’ code generation capabilities to develop critical software components (Suri et al., 2023). Numerous studies have explored GPTs’ ability to create web applications, knowledge graphs, and databases using prompt templates and fine-tuning techniques (Liukko et al., 2024; Jamdade and Liu, 2024; Vamsi et al., 2024). Since urban digital twins are essentially web-based or cloud-based applications, their core software components could potentially be generated using GPTs and other pre-trained language models. Consequently, we envision an autonomous system for code generation as a component of the service layer in CDTs, designed to automate the time-consuming software development and maintenance tasks. The approach can generate codes for both the server-side and client-side applications.
Integrating GenAI-powered autonomous systems (i.e., agents) into a CDT can significantly lower the technical barriers and programming skills required for developing an urban digital twin. In terms of new application opportunities, this paradigm can reduce the costs associated with software development and maintenance while increasing the efficiency of domain experts, such as urban planners and researchers in academia with moderate coding experience, allowing them to develop robust urban digital twin applications. These autonomous systems facilitate the creation of specialized research software, including digital twins, decision support systems, and content management systems. Consequently, researchers can focus more on their domain expertise and less on technical implementation and software engineering practices.This paradigm enable cost-effective solutions to address digital twin challenges C3 and C4 discussed in Section 1.2.
4.1.4 Human-AI Partnership
In the CDT’s cognition layer, where human involvements are often required to make critical decisions, GenAI models and their generative abilities can be harnessed to promote human-AI partnerships for a more participatory, adaptive, collaborative urban management approaches, which match the demanded and skills of various participants in urban managements. These participants include domain experts (e.g., urban planners and researcher), policymakers, and city residents. The proposed approaches are centered are around two aspects, namely conversational AI for decision support and GenAI-aided co-design.
Conversational AI
refers to emerging AI agents powered by advanced large language models, such as GPT, which enable intuitive and interactive communication between human users and AI systems (McTear, 2022). Unlike traditional expert systems that often require users to possess domain knowledge, conversational AI offers a more natural, intuitive, and user-friendly interface. This allows individuals to participate in decision-making processes informed by data-driven insights, domain-specific simulations, and AI-powered predictive analytics Sakirin and Said (2023). Conversational AI paradigms have proven effective in facilitating public engagement and participation in urban planning and management (Sahab et al., 2024). By providing a GPT-powered medium with excellent language understanding capabilities, these systems eliminate the need for users to have domain expertise. This approach enables interactive, user-centered dialogues with city residents and policymakers, capturing their needs and providing education and information relevant to support informed decisions. By integrating conversational AI into urban digital twins, GenAI promotes a more interactive, efficient, and democratic decision-making process, ultimately leading to more resilient and adaptive urban environments.
GenAI-aided Co-Design
can revolutionize urban planning and management by enabling more dynamic, collaborative, and innovative approaches to city development and management. GenAI models assist in co-design processes by interpreting and responding to user inputs, allowing stakeholders to collaboratively explore various planning scenarios and design alternatives. Leveraging the generative capabilities of models such as GANs, VAEs and generative diffusion models, trained on successful design blueprints and practices, domain experts (e.g., urban planners and researchers) can automate the creation of urban designs (e.g., floor plans, street layouts, landscapes, and parks) across large geographic regions in a short period of time (Guridi et al., 2024; Quan, 2022). This paradigm not only lowers costs but also enables urban planners to formulate design problems and explore more options through AI-Generated Content (AIGC) (Jiang et al., 2023). Furthermore, GenAI-powered tools facilitate co-design processes by allowing stakeholders, including community members, policymakers, and architects, to interact with AI systems through intuitive interfaces. This interaction can include generating and visualizing different design alternatives, optimizing layouts for sustainability and efficiency, and integrating inquiries and feedback in real-time (Knight et al., 2023).